After Your Run

** NOTE: These pages (BYR) have not yet been updated - APJ 091111 **
 After Your Run contains:
  Data analysis [OLD]
Data analysis [NEW]
Data analysis [Latest!!]
-Procedure, help files -List of variables in pipeline: See /import/chef2/susi/analysis/variables.doc -Data summary - searchable -Observing Log - searchable
Various Documents
WJT's document regarding the effective wavelength.
JD's  document regarding the effective wavelength.
The clocks on gareth and gaheris sometimes become out of sync leading to a 
time offset which affects the correlation between the wobble flux and fringe
scintillation. This is fixed by adding the keyword fixed_offset=?? to the
call to exrtract_wobble in check_v2_yymmdd.script.

The following fixed_offsets have been found so far:
 3601.5          040414, 040415
 3601.4          040416
 3601.3          040417
 3601.2          040418

 10.6            050107, 050108, 050109, 050110, 050111, 050112, 050114
                 050115, 050116

Dark counts
Up to and including 20/03/2004      -700nm filter, old APD0: 480\pm300 cts/s
                                     700nm filter, old APD1: 300\pm300 cts/s
21/03/2004 to 09/04/2004 inclusive  -700nm filter, new APDs: 800\pm500 cts/s
10/04/2004 onwards                  -700nm filter, new APDs: 20\pm10 cts/s
                                      with IR blocking filter
Changes to Observing practices
March 21,2004      - Installed new APDs, ie Red Table APDs aka 10504 & 10516
May 02, 2004       - Started using Scheduler
                     More regular foto files taken
May 12, 2004       - Started regular logging of pathcomp data but this didn't 
                     last long
                     pathcomp logs only in 040512, 040513, 040515, 040516,
                       040519, 040520, 040621(x3), 040628, 040629, 040701,
                       040718, 040719, 040727(x11)
~Before July 22,2004 - Started regularly taking wobble data, but in large
                       .wob files
July 22, 2004      - Foto file numbering still out of sync
Oct 02, 2004       - Foto file numbering now in sync
July 11,2005       - Started using separate .wob2 files. But the first data
                     with separate .wob2 files appears to be on 050717.
                     Wobble chopping should not be necessary on any data after
Nov 01, 2005       - Switched from SITE camera back to EEV camera for wobble2.
Dec 10, 2005	   - First observations with dichroics in. Dichroics installed
                      between 4 and 10 Dec.
                   - APD2 dead. Replaced briefly by APD1, before installing 
                     new blue APDs.
Dec 13, 2005       - Started using Blue Table APDs (11641 & 11638) on Red Table
Oct 01, 2006	   - JRN installed new with correct APD deadtime values.
Jan 11, 2007	   - APJ corrected extract_wobble. Changed type=1 error in call
                     to calc_centroids. See
Jan 19, 2007	   - APJ installed new pipeline at Sydney (1800hrs EASummerT)
Jan 20, 2007	   - APJ installed new pipeline at Narrabri (1330hrs EASummerT)
                     This 'new pipeline' corrects errors in the pipeline. See
                     Pipeline_corrections.pdf for details.
Feb 07, 2007	   - APJ installed new (@1800EASummerT in 
                     Sydney and Narrabri). This corrects for
                     noise spikes before doing correlation and before finding
                     dark frames. See WobFr_Correlation.doc 
                     (Sorry! Can't convert this one to a pdf) for details.
Feb 14, 2007       - APJ Fixed a bug in that caused trouble
                     in the dark frame search if fluxtot<0
		   - APJ Small cosmetic change to check_v2.script. Put ***s
                     in output between fringe files to improve readability.
May 15, 2007	   - School of Physics support switched to IDL 6.3 at 11:30am.
                     Before this we were using IDL 6.1
Aug 17, 2007	   - School of Physics support updated to IDL 6.4 between 8am
                      and 9am.

Andrew Jacob
Last modified: Fri May 28 14:27:13 EST 2004